LECTURE SCHEDULE, MAT 125, Spring 2004

Week of  Sections  Holidays and Exams

1/26  1.1, 1.2, 1.3
2/2  1.5, 1.6, 2.1
2/9 2.2, 2.3 Early Exam, Wed Feb 11, 8:30pm
2/16  2.4, review
2/23 2.5, 2.6 First Exam, Mon Feb 23, 8:30pm
3/1  2.7, 2.8, 2.9
3/8  2.10, 3.1
3/15  3.2, 3.4
3/22  3.5, 3.6
3/29 review, 3.7 Second Exam, Tue Mar 30, 8:30pm
4/5  No class Mon-Fri
4/12  4.1, 4.2
4/19  4.3, 4.5
4/26  4.6, 4.7
5/3 4.8, review Fri May 9 is last day of classes

5/10

5/17  Final, Mon May 17, 11-1:30

Rooms for the midterms and finals will be announced in class.

HOMEWORK, MAT 125, Spring 2004

Section: Topic  Problems:
1.1  Four ways to represent a function  2,18,19,36,38
1.2  Mathematical Models  4,9,11,
1.3  New functions from old  18,31,35,40,44,50
1.5  Exponential functions  10,15,17,23
1.6  Inverse functions  10,24,32,40,57
2.1  Tangent and velocity problems  4,7
2.2  Limit of a function  4,6,10,11,18
2.3  Calculating limits by limit laws  4,7,14,19,20,31
2.4  Continuity  4,14,27,31,32
2.5  Limits involving infinity  4,14,19,23,28,35,36,43
2.6  Tangents, velocities and rates of change 3,7,13,20
2.7  Derivatives  2,6,7,10a,16,21
2.8  Derivative as a function  3,14,22,25,36,47
2.9  Linear approximations  2,7,13,14
2.10  What does f′ say about f?  6,10,12,13,18,24,25
3.1  Polynomials and exponential functions 6,11,18,38,43,46,50
3.2  Product and quotient rules  6,11,16,20,28,31,36
3.4  Trigonometric functions  4,8,12,17,28,29
3.5  Chain rule  9,19,25,49,54,56,68
3.6  Implicit differentiation  5,7,10,17,23,30,54
3.7  Logarithmic functions  7,14,20,29,38
4.1  Related rates  7,9,12,14,17
4.2  Maximum and minimum values  6,24,29,40
4.3  Shapes of curves  6,7,17,28,48
4.5  l’Hospital’s rule  5,11,15,21,27,51
4.6  Optimization  2,5,10,11,13,14
4.8  Newton’s method  4,6,8,13,23